OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = x +

OpenStudy (anonymous):

medal and fan

OpenStudy (anonymous):

y2-y1 ----- x2-x1 is slope. can you fill that in?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

slope would be 9-1 ----- 2-3

OpenStudy (anonymous):

ohh i get u now.

OpenStudy (anonymous):

m = (y2 - y1)/(x2 - x1). m = (9 - 1)/(2 - 3). m = 8/-1 = - 8. Therefore slope, m = -8. Using y = mx + c. Slope = m, m = -8. y = -8x + c Substituting any of the point (3,1) into the equation, x = 3, y = 1. 1 = -8(3) + c 1 = -24 + c 1+24 = c 25 = c c = 25 Substituting into y = mx + c, m =-8, c = 25 Therefore y = -8x + 25. y = -8x + 25.

OpenStudy (anonymous):

thx sooo much.